Q . FOR THE FOLLOWING EQUILIBRIUM :-
 PCl5 (g) ---------- --------->PCl3(g) + Cl2(g)
VAPOUR DENSITY IS FOUND TO BE 100 WHEN 1 MOLE OF PCl5 IS TAKEN IN 10 LITRE FLASK AT 270C . THE PERCENTAGE DEGREE OF DISSOCIATION OF PCl5 IS-
  1. 5.5%
  2. 6.25%
  3. 2.12%
  4. 4.25%
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 

Grade:Select Grade

1 Answers

Sunil Kumar FP
9 years ago
ANS-(4)
If the equilibrium reaction is established in a closed vessel, then vapor density will be inversely proportional to the number of moles of the gaseous species as the density of the gaseous mixture (r) is a constant.

inital vapour density/final vapour density =number of moles at equilibrium/initial mole
we have
initial mole=1
initial vapour density= molecular formula of PCL5/2 =208.23/2 =104
final vapour density =100
PCL5----PCL3 +CL2
inital 1 0 0
final 1-d d d
number of moles at equilibrium =1+d,where d is the degree of dissociation
104/100 =1+d/1
1.04 =1+d
d=.04 or 4%